OpenStudy (anonymous):

I GIVE MEDALS AND FAN! Find the equation of the ellipse with the following properties. The ellipse with x-intercepts (5, 0) and (-5, 0); y-intercepts (0, 3) and (0, -3).

OpenStudy (anonymous):

centre (0,0) x^2/ a^2+y^2/9=1 a^2=16+9 =25 so a=5 Equation: x^2 /25 + y^2 /9 =1 http://www.wolframalpha.com/input/?i=x%5E2%2F25%2By%5E2%2F9%3D1